# Scaling the glassy dynamics of active particles: Tunable fragility and reentrance

### v1

@@ Abstract
Understanding the influence of activity on dense amorphous assemblies is crucial for biological processes such as wound healing, embryogenesis, or cancer progression. Here, we study the effect of self-propulsion forces of amplitude $f_{0}$ and persistence time $\tau_{p}$ in dense assemblies of soft repulsive particles, a model system that interpolates between particulate active matter and biological tissues. We identify the fluid and glass phases of the three-dimensional phase diagram obtained by varying $f_{0}$, $\tau_{p}$, and the packing fraction $\phi$. The morphology of the phase diagram directly accounts for a non-monotonic evolution of the relaxation time with $\tau_{p}$, which is a direct consequence of the crossover in the dominant relaxation mechanism, from glassy to jamming. A second major consequence is the evolution of the glassy dynamics from sub-Arrhenius to super-Arrhenius. We show that this tunable glass fragility extends to active systems analogous observations reported for passive particles. This allows us to apply a dynamic scaling analysis proposed for the passive case, in order to account for our results for active systems. Finally, we discuss similarities between our results and recent findings in the context of computational models of biological tissues.
@@ I Introduction
The effects of active processes on glassy dynamics have fundamental importance in several biological processes, such as wound healing Poujade et al. (2007); Vishwakarma et al. (2020), embryogenesis Mongera et al. (2018); Park et al. (2016); Atia et al. (2021); Tambe et al. (2011); Schötz et al. (2013), asthma development Park et al. (2015), or cancer progression Friedl and Wolf (2003); Friedl and Gilmour (2009); Malmi-Kakkada et al. (2018); Sinha et al. (2020); Palamidessi et al. (2019); Streitberger et al. (2020); Mitchel et al. (2020). The collective cellular dynamics during these processes exhibit a transition from a solid-like to a fluid-like state, although the static properties remain nearly the same Angelini et al. (2011); Berthier and Biroli (2011); Atia et al. (2021); Sadhukhan et al. (2024a); Berthier et al. (2019); Janssen (2019). The characteristics of the resulting glassy dynamics are broadly similar to those in particulate systems Berthier and Biroli (2011); Berthier et al. (2019); Sadhukhan et al. (2024a), although novel features may arise from the non-equilibrium nature of those systems. Indeed, cells are active and cellular systems are constantly evolving far from equilibrium Ramaswamy (2010); Marchetti et al. (2013). Cells can also change their characteristics during various processes Thiery (2002); Thiery et al. (2009), and they are extended objects for which cellular shapes closely correlate with biological functions Park et al. (2015); Sailem and Bakal (2017); Atia et al. (2018); Poujade et al. (2007); Chen et al. (1997). Geometric features can also differ from passive systems for the same reason; epithelial systems, for example, are always confluent, so that the cells entirely cover the space and the packing fraction may not be the most relevant parameter Atia et al. (2021).
@@
The complexity of active biological systems makes it challenging to gain insights into the key mechanisms driving their glassy dynamics. Simulation studies of model cellular systems have been instrumental here, showing a rigidity transition akin to the jamming transition Bi et al. (2015, 2016); Park et al. (2015). Furthermore, several studies have demonstrated that these systems exhibit nontrivial glassy behavior, such as sub-Arrhenius relaxation dynamics Sussman et al. (2018); Li et al. (2021); Sadhukhan and Nandi (2021), possibly crossing over to super-Arrhenius behavior as the shape index (which controls cell shape) varies Sadhukhan and Nandi (2021); Li et al. (2021); Sadhukhan et al. (2024b). It remains unclear whether these features are specific to biological tissues or can be observed in active particles, or even in equilibrium systems.
@@
Active systems composed of self-propelled particles do not capture all microscopic details of biological tissues, but may nevertheless provide useful insight into the competition between crowding and activity. In these models, particles self-propel under the influence of self-propulsion forces of amplitude $f_{0}$ and persistence time $\tau_{p}$ Ramaswamy (2010); Marchetti et al. (2013); Sadhukhan et al. (2024a); Berthier et al. (2019); Janssen (2019). In fact, several biological systems can be conveniently modelled as dense systems of self-propelled particles on different lengthscales, such as for example collections of cells Angelini et al. (2011); Garcia et al. (2015), the cytoplasm Zhou et al. (2009); Parry et al. (2014); Nishizawa et al. (2017), aggregates of organisms Gravish et al. (2013); Lama et al. (2024), and synthetic active matter Deseigne et al. (2010); Arora et al. (2022); Palacci et al. (2013); Lam et al. (2015). Several experiments Fabry et al. (2001); Angelini et al. (2011); Lenormand et al. (2007); Arora et al. (2022) and simulation studies Berthier (2014); Mandal et al. (2016, 2017); Flenner et al. (2016); Berthier et al. (2017) have shown that active systems in their dense regime exhibit many characteristics similar to glasses Berthier et al. (2019); Janssen (2019); Atia et al. (2021); Sadhukhan et al. (2024a), such as the anomalously slow two-step relaxation dynamics Park et al. (2015) and dynamical heterogeneity Berthier et al. (2011). Recent studies have also revealed nontrivial effects of activity on the glassy dynamics, such as a shifted glass transition point Berthier and Kurchan (2013); Berthier (2014); Nandi and Gov (2017), changing fragility Mandal et al. (2016); Flenner et al. (2016); Nandi et al. (2018), and re-entrant relaxation dynamics Berthier et al. (2017); Flenner et al. (2016); Debets et al. (2021). These active particle systems thus provide a rich yet still relatively simple framework for understanding the role of nonequilibrium fluctuations on the glassy dynamics.
@@
Theories of the equilibrium glass transition, such as the mode-coupling theory Berthier and Kurchan (2013); Szamel et al. (2015); Szamel (2016); Feng and Hou (2017); Liluashvili et al. (2017); Nandi and Gov (2017); Pandey et al. (2025); Paul et al. (2023); Kolya et al. (2024) and the random first-order transition approach Nandi et al. (2018); Mandal et al. (2022); Sadhukhan and Nandi (2021); Sadhukhan et al. (2024b) have been extended to active glasses to rationalize numerical and experimental results. In fact, several aspects of activity-driven glassy dynamics remain equilibrium-like at a suitably defined effective temperature Mandal et al. (2016); Nandi et al. (2018). However, there are also a number of nontrivial aspects that have been reported, such as re-entrant glassy dynamics Berthier et al. (2017); Debets et al. (2021); Arora et al. (2022), change of glass fragility Mandal et al. (2016); Flenner et al. (2016); Nandi et al. (2018), and emerging velocity correlations Szamel et al. (2015). Furthermore, the specific role of $\tau_{p}$ on the glassy dynamics has not been elucidated fully. Crucially, how to extend and connect these results to studies of more complex biological systems remains unclear.
@@
To address these issues and connect the glassy dynamics of particulate matter on the one hand and dense cellular biological tissues on the other, we investigate the active glassy dynamics in a system of soft repulsive particles. This model smoothly connects and unifies distinct physics in appropriate limits. In the limit of zero activity, the finite-temperature ($T$) glassy behavior is governed by a “glass point” in the limit of small but nonzero $T$, and a jamming transition at $T=0$ itself. One can thus explore both jamming and glassy physics within the same system Berthier and Witten (2009a, b); Berthier et al. (2010); Jacquin et al. (2011); Adhikari et al. (2023). On the other hand, the same system in the very dense limit is a good approximation of confluent cellular systems. Between these two limits the system describes the active matter physics of particulate systems, where qualitatively different regimes can be expected as the persistence time is varied. Therefore, this is a fairly rich model, which as we will demonstrate exhibits several non-trivial features that have also been reported in biological tissues. Our goal is thus to provide a physical interpretation of these observed behaviours in the context of a model that is easier to study, hopefully shedding light on biological systems.
@@ II Model for soft active particles
We consider a three-dimensional 50:50 binary mixture of particles of two types A and B, interacting via the Weeks-Chandler-Andersen purely repulsive potential. We assume that thermal fluctuations are not relevant and set the Brownian temperature to $T=0$. Instead, the particles are driven by non-thermal self-propulsion forces, for which we use the AOUP (active Ornstein-Uhlenbeck particle) activity model (see supplementary material (SM), Sec. S1 for further details). We use molecular dynamics simulations to evolve the particle position $\mathbf{r}_{i}$ of the $i$th particle using the following equation of motion:
@@
where $\mathbf{F}_{i}$ is the interaction force felt by the $i$th particle from the other particles and $\mathbf{f}_{i}$ is the active self-propulsion force. The latter follows an Ornstein-Uhlenbeck stochastic process
$\tau_{p}\dot{\mathbf{f}}_{i}=-\mathbf{f}_{i}+\bm{\eta}_{i},$ (2)
with $\langle\bm{\eta}_{i}(t)\bm{\eta}^{\rm T}_{j}(t^{\prime})\rangle=2\xi_{0}f_{0}^{2}\mathbf{I}\delta_{ij}\delta(t-t^{\prime})$, where “T” denotes the transpose, $\mathbf{I}$ the identity matrix and $\xi_{0}$ the friction coefficient. We present results using $\xi_{0}\sigma^{2}_{BB}/\epsilon$ as the time unit, with $\sigma_{BB}$ the diameter of B-type particles that sets the unit length, and $\epsilon$ the energy scale of the pair potential.
The self-propulsion force $\mathbf{f}_{i}$ in Eq. (S2) is stochastic with zero mean and correlator $\langle\mathbf{f}_{i}(t)\mathbf{f}^{\rm T}_{j}(0)\rangle=(f_{0}^{2}/\tau_{p})\exp[-t/\tau_{p}]\mathbf{I}\delta_{ij}$. It thus has a typical amplitude $|\mathbf{f}_{i}|\sim f_{0}/\sqrt{\tau_{p}}$ and fluctuations that are correlated over a time of order $\tau_{p}$. In the limit of very small persistent times, $\tau_{p}\to 0$, the propulsion forces thus become equivalent to a thermal white noise, with $f_{0}^{2}$ playing the role of a temperature. In the opposite limit of large persistence times, the propulsions become nearly constant random driving forces of amplitude $f_{0}/\sqrt{\tau_{p}}$.
@@
where $G$ is a constant Nandi and Gov (2017). This effective temperature reduces to $T_{\text{eff}}\sim f_{0}^{2}$ for small persistence times, and to $T_{\text{eff}}\sim f^{2}_{0}/\tau_{p}$ for larger $\tau_{p}$. The expression in Eq. (3) can be analytically justified by considering the position fluctuations of a single AOUP in a harmonic potential Szamel (2014). Using the effective temperature description for the relaxation dynamics and comparing it with the mode-coupling theory scaling relations (see SM Fig. S5), we estimate $G\simeq 0.5$ for the current system.
@@
The three control parameters of the model are the persistence time $\tau_{p}$, the amplitude $f_{0}$ in the noise term in Eq. (S3) and the volume fraction $\phi$. We will systematically vary them to fully characterise the part of the phase diagram where the system becomes glassy. It is first interesting to consider the various physical limits captured by this model.
@@
Small persistence time, $\tau_{p}\to 0$. The phase diagram becomes effectively two-dimensional in the equilibrium limit $\tau_{p}\to 0$, where the system becomes equivalent to a Brownian system with the two control parameters temperature and density, ($T_{\text{eff}}\sim f_{0}^{2}$, $\phi$), as usual in equilibrium fluids. The smooth approach to the Brownian limit makes the AOUP model in Eqs. (S2, S3) appealing. In this limit, the physics becomes strictly equivalent to the one of soft repulsive spheres at thermal equilibrium, as studied extensively in Ref. Berthier and Witten (2009b).
@@
Small force amplitude, $f_{0}\to 0$. For finite $\tau_{p}$, another interesting limit is the one where $f_{0}\to 0$. In this limit, the amplitude of the self-propulsion force $\mathbf{f}_{i}$ in Eq. (S2) becomes much smaller than that of the interaction forces $\mathbf{F}_{i}$, so that particles will no longer interpenetrate (unless forced to do so by a high volume fraction $\phi$). As a consequence, the system behaves as self-propelled hard spheres with a finite persistence time $\tau_{p}$. The phase diagram is again two-dimensional, with $(\tau_{p},\phi)$ as control parameters. Such self-propelled hard spheres have been studied in several past works Ni et al. (2013); Berthier (2014) and undergo a non-equilibrium glass transition, at a volume fraction $\phi_{d}(\tau_{p})$ that depends continuously on the persistence time. These special “glass points” in the phase diagram will play a prominent role in our dynamic scaling analysis in Sec. VII.
@@
Large persistence time, $\tau_{p}\to\infty$. The final interesing limit is that of infinitely persistent particles obtained when $\tau_{p}\to\infty$. In that limit, particles are driven by frozen random forces whose amplitudes $|f_{i}|\sim f_{0}/\sqrt{\tau_{p}}$ become vanishingly small compared to interparticle forces, in the absence of any other source of fluctuations. By construction, the system undergoes in this limit a jamming transition at a packing fraction $\phi_{J}$, between a flowing fluid and a solid phase. We expect the dynamics in this “jamming” limit to differ qualitatively from the glassy dynamics observed near glass transitions Ikeda et al. (2012).

### v2

# Scaling the glassy dynamics of active particles: Tunable fragility and re-entrance

@@ Abstract
Understanding the influence of activity on dense amorphous assemblies is crucial for biological processes such as wound healing, embryogenesis, or cancer progression. Here, we study the effect of self-propulsion forces of amplitude $f_{0}$ and persistence time $\tau_{p}$ in dense assemblies of soft repulsive particles by simulating a model particle system that interpolates between particulate active matter and biological tissues. We identify the fluid and glass phases of the three-dimensional phase diagram obtained by varying $f_{0}$, $\tau_{p}$, and the packing fraction $\phi$. The morphology of the phase diagram directly accounts for a non-monotonic evolution of the relaxation time with $\tau_{p}$, which is a direct consequence of the crossover in the dominant relaxation mechanism, from glassy to jamming. A second major consequence is the evolution of the glassy dynamics from sub-Arrhenius to super-Arrhenius. We show that this tunable glass fragility extends to active systems analogous observations reported for passive particles. This analogy allows us to apply a dynamic scaling analysis proposed for the passive case, in order to account for our results for active systems. Finally, we discuss similarities and differences between our results and recent findings in the context of computational models of biological tissues.
@@ I Introduction
The effects of active processes on glassy dynamics have fundamental importance in several biological processes, such as wound healing Poujade et al. (2007); Vishwakarma et al. (2020), embryogenesis Mongera et al. (2018); Park et al. (2016); Atia et al. (2021); Tambe et al. (2011); Schötz et al. (2013), asthma development Park et al. (2015), or cancer progression Friedl and Wolf (2003); Friedl and Gilmour (2009); Malmi-Kakkada et al. (2018); Sinha et al. (2020); Palamidessi et al. (2019); Streitberger et al. (2020); Mitchel et al. (2020). The collective cellular dynamics during these processes exhibit a transition from a solid-like to a fluid-like state, although the static properties remain nearly the same Angelini et al. (2011); Berthier and Biroli (2011); Atia et al. (2021); Sadhukhan et al. (2024a); Berthier et al. (2019); Janssen (2019). The characteristics of the resulting glassy dynamics are broadly similar to those in equilibrium particulate systems Berthier and Biroli (2011); Berthier et al. (2019); Sadhukhan et al. (2024a); Higler et al. (2018), although novel features may arise from the non-equilibrium nature of the systems. Indeed, cells are active and cellular systems are constantly evolving far from equilibrium Ramaswamy (2010); Marchetti et al. (2013a). Cells can also change their characteristics during various processes Thiery (2002); Thiery et al. (2009), and they are extended objects for which cellular shapes closely correlate with biological functions Park et al. (2015); Sailem and Bakal (2017); Atia et al. (2018); Poujade et al. (2007); Chen et al. (1997). Geometric features can also differ from passive systems for the same reason; epithelial systems, for example, are always confluent, so that the cells entirely cover the space and the packing fraction may not be the most relevant parameter Atia et al. (2021).
@@
The complexity of active biological systems makes it challenging to gain insights into the key mechanisms driving their glassy dynamics. Simulation studies of model cellular systems have been instrumental here, showing a rigidity transition akin to the jamming transition Bi et al. (2015, 2016); Park et al. (2015). Furthermore, several computational studies have demonstrated that these systems exhibit nontrivial glassy behavior, such as sub-Arrhenius relaxation dynamics Sussman et al. (2018); Li et al. (2021); Sadhukhan and Nandi (2021), possibly crossing over to super-Arrhenius behavior as the shape index (which controls cell shape) varies Sadhukhan and Nandi (2021); Li et al. (2021); Sadhukhan et al. (2024b). It would be useful to elucidate whether these features are specific to biological tissues and vertex models, or can be observed in simpler models of active particles.
@@
Active systems composed of self-propelled particles cannot capture all microscopic details of biological tissues, but may nevertheless provide useful insights into the competition between crowding and activity. Beyond this they can also reveal what features are specific to confluent models with many-body interactions, compared to particle models with pairwise forces. In the latter models, particles self-propel under the influence of self-propulsion forces of amplitude $f_{0}$ and persistence time $\tau_{p}$ Ramaswamy (2010); Marchetti et al. (2013a); Sadhukhan et al. (2024a); Berthier et al. (2019); Janssen (2019); Mandal et al. (2020); Mandal and Sollich (2020); Wiese et al. (2023); Szamel and Flenner (2024). In fact, several biological systems can be conveniently modelled as dense systems of self-propelled particles on different lengthscales, such as for example collections of cells Angelini et al. (2011); Garcia et al. (2015), the cytoplasm Zhou et al. (2009); Parry et al. (2014); Nishizawa et al. (2017), aggregates of organisms Gravish et al. (2013); Lama et al. (2024), and synthetic active matter Deseigne et al. (2010); Arora et al. (2022); Palacci et al. (2013); Lam et al. (2015). Several experiments Fabry et al. (2001); Angelini et al. (2011); Lenormand et al. (2007); Arora et al. (2022) and simulation studies Berthier (2014); Mandal et al. (2016, 2017); Flenner et al. (2016); Berthier et al. (2017) have shown that active systems in their dense regime exhibit many characteristics similar to glasses Berthier et al. (2019); Janssen (2019); Atia et al. (2021); Sadhukhan et al. (2024a), such as the anomalously slow two-step relaxation dynamics Park et al. (2015) and dynamical heterogeneity Berthier et al. (2011). Recent numerical studies have also revealed nontrivial effects of activity on the glassy dynamics, such as a shifted glass transition point Berthier and Kurchan (2013); Berthier (2014); Nandi and Gov (2017), changing fragility Mandal et al. (2016); Flenner et al. (2016); Nandi et al. (2018); Dey and Karmakar (2025), and re-entrant relaxation dynamics Berthier et al. (2017); Flenner et al. (2016); Debets et al. (2021a). These active particle systems thus provide a rich yet still relatively simple framework for understanding the role of nonequilibrium fluctuations on the glassy dynamics.
There are several variants of active matter models in the literature Marchetti et al. (2013b), sometimes categorized into dry and wet active matter. In the former, the surrounding fluid is either absent or treated implicitly within the models. In the latter, the surrounding fluid is explicitly taken into account. Since glassy effects arise in the long-time dynamics, we expect the hydrodynamic role of the surrounding fluid to be insignificant, as it affects the physics on much shorter timescales. We thus believe that the dry-wet distinction is less relevant in the present context Klongvessa et al. (2019, 2022). Similarly, dimensionality effects are not expected to play an important role for the specific questions addressed in our study.
@@
Theories of the equilibrium glass transition, such as the mode-coupling theory Berthier and Kurchan (2013); Szamel et al. (2015); Szamel (2016); Feng and Hou (2017); Liluashvili et al. (2017); Nandi and Gov (2017); Paul et al. (2023); Debets et al. (2021b); Paul et al. (2023); Debets and Janssen (2023); Kolya et al. (2024); Pandey et al. (2025) and the random first-order transition approach Nandi et al. (2018); Mandal et al. (2022); Sadhukhan and Nandi (2021); Sadhukhan et al. (2024b) have been extended to active glasses to rationalize numerical and experimental results. In fact, several aspects of activity-driven glassy dynamics remain equilibrium-like at a suitably defined effective temperature Mandal et al. (2016); Nandi et al. (2018). However, there are also a number of nontrivial aspects that have been reported, such as re-entrant glassy dynamics Berthier et al. (2017); Debets et al. (2021a); Arora et al. (2022), change of glass fragility Mandal et al. (2016); Flenner et al. (2016); Nandi et al. (2018); Dey and Karmakar (2025), and emerging velocity correlations Szamel et al. (2015). Furthermore, the specific role of $\tau_{p}$ on the glassy dynamics has not been elucidated fully. Crucially, how to extend and connect these results to studies of more complex biological systems remains unclear.
@@
To address all these issues, we investigate active glassy dynamics via computer simulations of a system of soft repulsive particles. This model smoothly connects and unifies distinct physics in appropriate limits. In the limit of zero activity, the finite-temperature ($T$) glassy behavior is governed by a “glass point” in the limit of small but nonzero $T$, and a jamming transition at $T=0$ itself. One can thus explore both jamming and glassy physics within the same system Berthier and Witten (2009a, b); Berthier et al. (2010); Jacquin et al. (2011); Adhikari et al. (2023). On the other hand, the same system in the very dense limit is a good approximation of confluent cellular systems. Between these two limits the system describes the active matter physics of particulate systems, where qualitatively different regimes can be expected as the persistence time is varied. Therefore, this is a fairly rich model, which as we will demonstrate exhibits several non-trivial features that have also been reported in biological tissues. Our goal is, then, to provide a physical interpretation of the observed behaviors, in the context of a particle model that is relatively easy to study and understand. This approach will also be useful to guide and interpret future studies of biological materials and tissue models, for instance to disentangle specific effects due to the confluent nature of tissues or the possible influence of many-body forces.
@@ II Model for soft active particles
We consider a three-dimensional 50:50 binary mixture of particles of two types A and B, interacting via the Weeks-Chandler-Andersen purely repulsive potential. We expect that similar results will hold in two dimensions, as the effects of dimensionality on glassy dynamics are well understood Flenner and Szamel (2015). We assume that thermal fluctuations are not relevant and set the Brownian temperature to $T=0$. Instead, the particles are driven by non-thermal self-propulsion forces, for which we use the AOUP (active Ornstein-Uhlenbeck particle) activity model (see supplementary material (SM), Sec. S1 for further details). Previous studies have demonstrated that similar physics would be obtained independently of the specific model chosen for the self-propulsion or the repulsive pairwise interaction Debets and Janssen (2022).
We use molecular dynamics simulations to evolve the particle position $\mathbf{r}_{i}$ of the $i$th particle using the following equation of motion:
@@
where $\xi_{0}$ the friction coefficient, $\mathbf{F}_{i}$ is the interaction force felt by the $i$th particle from the other particles, and $\mathbf{f}_{i}$ is the active self-propulsion force. The latter follows an Ornstein-Uhlenbeck stochastic process
$\tau_{p}\dot{\mathbf{f}}_{i}=-\mathbf{f}_{i}+\boldsymbol{\eta}_{i},$ (2)
with $\langle\boldsymbol{\eta}_{i}(t)\boldsymbol{\eta}^{\rm T}_{j}(t^{\prime})\rangle=2f_{0}^{2}\mathbf{I}\delta_{ij}\delta(t-t^{\prime})$, where “T” denotes the transpose and $\mathbf{I}$ the identity matrix. We present results using $\xi_{0}\sigma^{2}_{BB}/\epsilon$ as the time unit, with $\sigma_{BB}$ the diameter of B-type particles that sets the unit length, and $\epsilon$ the energy scale of the pair potential.
The self-propulsion force $\mathbf{f}_{i}$ in Eq. (1) is stochastic with zero mean and correlator $\langle\mathbf{f}_{i}(t)\mathbf{f}^{\rm T}_{j}(0)\rangle=(f_{0}^{2}/\tau_{p})\exp[-t/\tau_{p}]\mathbf{I}\delta_{ij}$. It thus has a typical amplitude $|\mathbf{f}_{i}|\sim f_{0}/\sqrt{\tau_{p}}$ and fluctuations that are correlated over a time of order $\tau_{p}$. In the limit of very small persistent times, $\tau_{p}\to 0$, the propulsion forces thus become equivalent to a thermal white noise, with $f_{0}^{2}$ playing the role of a temperature. In the opposite limit of large persistence times, the propulsions become nearly constant random driving forces of amplitude $f_{0}/\sqrt{\tau_{p}}$.
@@
where $G$ is a constant Nandi and Gov (2017). This effective temperature reduces to $T_{\text{eff}}\sim f_{0}^{2}$ for small persistence times, and to $T_{\text{eff}}\sim f^{2}_{0}/\tau_{p}$ for larger $\tau_{p}$. The expression in Eq. (3) can be analytically justified by considering the position fluctuations of a single AOUP in a harmonic potential Szamel (2014).
The concept of an effective temperature has a long history in both active and glass matter Cugliandolo (2011). For active glasses, an effective temperature naturally emerges in the long-time dynamics Berthier and Kurchan (2013); Levis and Berthier (2015); Nandi and Gov (2017). It has even been shown that the relaxation dynamics for the active system agrees well with the mode-coupling scaling description using $T_{\text{eff}}$. Using this effective temperature description for the relaxation dynamics and comparing it with the mode-coupling theory scaling relations (see SM Fig. S5), we estimate $G\simeq 0.5$ for the current system. For the present purposes, $T_{\text{eff}}$ is a useful quantity as it reduces to a genuine thermodynamic temperature in the small $\tau_{p}$ limit, but our results do not rely on any thermodynamic interpretation far from equilibrium.
@@
The three control parameters of the model are the persistence time $\tau_{p}$, the amplitude $f_{0}$ in the noise term in Eq. (2) and the volume fraction $\phi$. We will systematically vary them to fully characterise the part of the phase diagram where the system becomes glassy. It is first interesting to consider the various physical limits captured by this model.
@@
Small persistence time, $\tau_{p}\to 0$. The phase diagram becomes effectively two-dimensional in the equilibrium limit $\tau_{p}\to 0$, where the system becomes equivalent to a Brownian system with the two control parameters temperature and density, ($T_{\text{eff}}\sim f_{0}^{2}$, $\phi$), as usual in equilibrium fluids. The smooth approach to the Brownian limit makes the AOUP model in Eqs. (1, 2) appealing. In this limit, the physics becomes strictly equivalent to the one of soft repulsive spheres at thermal equilibrium, as studied extensively in Ref. Berthier and Witten (2009b).
@@
Small force amplitude, $f_{0}\to 0$. For finite $\tau_{p}$, another interesting limit is the one where $f_{0}\to 0$. In this limit, the amplitude of the self-propulsion force $\mathbf{f}_{i}$ in Eq. (1) becomes much smaller than that of the interaction forces $\mathbf{F}_{i}$, so that particles will no longer interpenetrate (unless forced to do so by a high volume fraction $\phi$). As a consequence, the system behaves as self-propelled hard spheres with a finite persistence time $\tau_{p}$. The phase diagram is again two-dimensional, with $(\tau_{p},\phi)$ as control parameters. Such self-propelled hard spheres have been studied in several past works Ni et al. (2013); Berthier (2014) and undergo a non-equilibrium glass transition, at a volume fraction $\phi_{d}(\tau_{p})$ that depends continuously on the persistence time. These special “glass points” in the phase diagram will play a prominent role in our dynamic scaling analysis in Sec. VII.
@@
Large persistence time, $\tau_{p}\to\infty$. The final interesting limit is that of infinitely persistent particles obtained when $\tau_{p}\to\infty$. In that limit, particles are driven by frozen random forces whose amplitudes $|f_{i}|\sim f_{0}/\sqrt{\tau_{p}}$ become vanishingly small compared to interparticle forces, in the absence of any other source of fluctuations. By construction, the system undergoes in this limit a jamming transition at a packing fraction $\phi_{J}$, between a flowing fluid and a solid phase. We expect the dynamics in this “jamming” limit to differ qualitatively from the glassy dynamics observed near glass transitions Ikeda et al. (2012).